### v2

[…]
1 Introduction
Modern large-scale language models (Brown et al. 2020, Chowdhery et al. 2023, OpenAI 2023, Kaplan et al. 2020, Wei et al. 2022) have revolutionized artificial intelligence by demonstrating unprecedented capabilities in natural language generation across diverse linguistic domains and situational contexts. These sophisticated neural networks, trained on extensive corpora of textual data, now serve as foundational components for numerous real-world applications. Their deployment spans conversational agents (Anthropic 2023, Character 2021, OpenAI 2019, 2023), information retrieval systems (Microsoft 2023, Google 2023, Komo 2023, Perplexity 2022, You.com 2020), programming aids (Amazon 2023, GitHub 2021, Replit 2018), and even clinical decision support tools (Cascella et al. 2023, Peng et al. 2023, Sallam 2023), demonstrating remarkable versatility in professional and consumer domains alike.
Efficient inference represents a fundamental challenge in large language model deployment, focusing on how models process input prompts and generate responses with minimal latency. The inference pipeline consists of two distinct phases. (1) Prefill: processing the input prompt (a user-submitted text sequence), and (2) Decode: autoregressively generating output tokens (where each token typically corresponds to a word or sub-word unit). Modern LLMs universally employ the Transformer architecture (Vaswani et al. 2017) for this process. For instance, when processing the prompt “Why is the sea blue?”, the model first tokenizes it into discrete units (“Why”, “is”, “the”, “sea”, “blue”, “?”), then sequentially generates output tokens (e.g., beginning with “Because”) while considering both the prompt and previously generated tokens at each step. The core computational challenge emerges in multi-request scenarios, where the system must optimally schedule numerous concurrent inference tasks. This scheduling problem requires careful balancing of computational resources to minimize user-perceived latency, particularly given the autoregressive nature of token generation and the variable length of both input prompts and output responses.
The scheduling problem in LLM inference presents unique challenges that distinguish it from traditional scheduling tasks, primarily due to its dynamic memory constraints. During inference, each computational worker maintains a fixed Key-Value (KV) cache memory capacity, where keys and values represent contextual embeddings derived from both input tokens and autoregressively generated output tokens. The memory capacity depends on the hardware of each worker and the complexity of the LLM used. This memory system exhibits two critical behaviors: (1) it must retain all processed tokens from the prefill input prompt, and (2) its consumption grows linearly during decode generation as each newly generated output token requires additional KV cache allocation. The memory growth pattern creates complex scheduling dynamics when combined with parallel processing requirements. Unlike traditional systems where task resource demands remain static, LLM inference exhibits variable memory pressure - the number of concurrent requests a worker can handle fluctuates dynamically based on each request’s generation progress. This nonlinear relationship between sequence length and memory consumption fundamentally transforms the scheduling task. The challenge is further compounded by the need to balance immediate memory availability against anticipated future demands as generation progresses across multiple concurrent requests.
Jaillet et al. (2025) established a theoretical model for LLM inference scheduling on a single computational worker, proposing a shortest-first scheduling policy that prioritizes requests with fewer output tokens. Their approach dynamically batches requests while respecting memory constraints, demonstrating both theoretical and numerical effectiveness under the condition of uniform input sizes. However, this condition may not always hold in practice, where workers typically handle mixed workloads (e.g., combining short conversational prompts with lengthy document summarization tasks).
Relaxing this condition fundamentally changes the scheduling problem. With variable prefill input sizes, the relationship between processing time, memory usage, and optimal scheduling becomes significantly more complex. For instance, a request with a large input but small output may consume substantial memory yet finish quickly, while one with a small input but long output may occupy a small memory for an extended period. This trade-off makes prioritization non-trivial, as the previously effective shortest-first heuristic becomes inadequate. In Section 2.1, we formally analyze these challenges and demonstrate that the generalized scheduling problem is NP-hard and that existing algorithms perform poorly without the uniform input size assumption.
1.1 Main Contribution
In this section, we introduce the main contributions and the roadmap of this paper.
Negative Results for Existing Scheduling Algorithms and NP-hardness. In Section 2, we first establish fundamental limitations of current methods by analyzing the model in Jaillet et al. (2025) under realistic conditions. Our key theoretical findings reveal that: (1) without uniform input sizes, both first-come-first-serve and shortest-first scheduling yield unbounded competitive ratios (regardless of using output length or total sequence length as the metric), and (2) the general scheduling problem is NP-hard. These results formally justify the challenge need for new scheduling approaches.
Polynomial Time Scheduling Algorithm with Constant Competitive Ratio. In Section 3, we introduce Sorted-F, a novel scheduling algorithm that achieves both computational efficiency and provable performance guarantees. The algorithm operates through three key steps: (1) it first evaluates each request using our designed quality metric, which comprehensively considers both input and output characteristics; (2) it then partitions requests into prioritized batches based on this metric; and (3) within each batch, it applies shortest-output-first prioritization. Crucially, we prove that Sorted-F maintains a constant competitive ratio of at most 48, independent of problem scale – a significant improvement over existing approaches that suffer from unbounded ratios in general cases. The proof, presented in Section 4, introduces novel techniques for handling variable input/output sizes, which may be of independent interest and inspire new theoretical analyses for similar scheduling problems.
Approximation Algorithms for Practical Deployment. Since scheduling latency directly impacts overall inference time in production systems, we investigate efficient approximations of Sorted-F that maintain its performance benefits while reducing computational overhead. The primary complexity bottleneck in Sorted-F lies in its batch partitioning phase based on our quality metric. In Section 5, we analyze four approximation approaches: (1) exact dynamic programming, (2) scaled dynamic programming, (3) local swap search, and (4) quantile greedy selection. For each method, we characterize both its theoretical computational complexity and practical applicability, specifying the problem scales where each variant proves to be most effective. These analyses provide system designers with clear guidelines for algorithm selection based on their specific latency requirements and workload characteristics.
LP-based Heuristics. In Section 6, we further analyze the scheduling problem through an optimization lens by formulating it as an Integer Program (IP), as suggested in Jaillet et al. (2025). While this IP provides theoretical optimality, its time complexity makes it impractical for real-time inference systems requiring millisecond-level decisions. Our work makes two key contributions in this direction: first, we discuss the gap between the IP formulation and its LP relaxation; second, we develop Sorted-LP, a novel heuristic that leverages the relaxed LP solution structure. Sorted-LP operates by (1) solving the LP relaxation, (2) extracting expected starting times from the fractional solution, and (3) sorting requests according to the expected starting times. We also characterize Sorted-LP’s advantages and limitations, providing practitioners with clear guidance for algorithm selection.
Numerical Experiments. In Section 7, we present comprehensive empirical evaluations using real-world data to validate our theoretical findings. To accurately capture the practical challenges of variable-length inputs, we construct a mixed dataset combining: (1) short conversational prompts from Zheng et al. (2023), and (2) long document summarization tasks from Cohan et al. (2018). This experimental design reflects realistic deployment scenarios where LLM servers must handle diverse request types simultaneously. We evaluate four scheduling approaches: first-come-first-serve, shortest-first, our LP-based Sorted-LP, and our main proposed algorithm Sorted-F. The results demonstrate that Sorted-F achieves significantly lower average latency, confirming both the theoretical robustness and practical effectiveness of our approach.
1.2 Other Related Works
LLM Inference. Improving the efficiency of LLM inference is crucial due to the substantial computational and financial costs it incurs. A large body of existing work focuses on practical implementation techniques to accelerate inference in deployed systems, often without accompanying theoretical guarantees. For instance, Patel et al. (2023), Zhong et al. (2024) propose system architectures that process prompt and token requests independently, while other works such as Yu et al. (2022), Agrawal et al. (2023, 2024b) introduce designs that jointly handle prompt and token requests—a structure aligned with the one studied in this paper.
In contrast, a few recent studies aim to develop theoretical foundations for LLM inference. The primary mathematical model considered in this paper builds on the framework introduced by Jaillet et al. (2025). Ao et al. (2025) examine inference scheduling under multiple objectives and propose algorithms with provable performance guarantees. Additionally, Li et al. (2025) analyze the stability conditions of LLM inference when operating on a single computational worker.
Scheduling. The scheduling problem—extensively studied in the literature (Allahverdi et al. 2008, Chen et al. 1998, Brucker et al. 1999, Albers 2009, Mak et al. 2015, Xing and Zhang 2000, Kong et al. 2013)—involves a decision-maker tasked with processing a large number of incoming requests, either one-by-one or in batches. The challenge lies in determining the optimal order and timing of processing, which is often formulated as an integer optimization problem. However, due to the computational complexity and the impracticality of solving large-scale instances exactly, researchers have developed fast approximation algorithms and heuristics.
In our setting, which is motivated by LLM inference, jobs can be processed in batches. This aligns our problem more closely with batch scheduling models studied in (Brucker et al. 1998, Chen and Lee 2008, Liu and Lu 2015, Li et al. 2020). Furthermore, LLM inference imposes a sequential structure on token generation—future tokens cannot be processed before current ones—introducing a form of precedence constraint. Related work on scheduling with precedence constraints includes (Shahout et al. 2024, Garg et al. 2019, Azar and Epstein 2002, Robert and Schabanel 2008, Agrawal et al. 2016). Nevertheless, our setting is fundamentally different due to the unique characteristics of the KV cache, which introduces new constraints and tradeoffs not captured by traditional precedence scheduling models.
2 Model Review
In this section, we review the model introduced by Jaillet et al. (2025). The setting considers LLM inference on a single computational worker with a key-value (KV) cache limit $M>0$, meaning that the worker can hold at most $M$ tokens in memory at any given time. In practice, $M$ depends on the size of the model and the hardware specifications (e.g., GPU memory), and is assumed to be known to the decision-maker.
We consider a discrete-time system where $n$ prompts arrive simultaneously at time $t=0$. Each prompt $i\in[n]$ is characterized by a pair $(s_{i},o_{i})$, where $s_{i}$ denotes the input size (i.e., the number of tokens in the input prompt), and $o_{i}$ denotes the number of output tokens to be generated. We assume that $o_{i}$ can be accurately predicted, and that both scale as $s_{i},o_{i}=\epsilon(M)$, where $\epsilon(M)$ is a term asymptotically smaller than $M$. This captures the natural regime where individual requests consume a vanishing fraction of total memory.
Each request $i$ must be processed token by token in order. The memory required to generate the $j$th token (where $j\in[o_{i}]$) is $s_{i}+j$, reflecting the growing size of the KV cache as new tokens are appended. At each time step $t$, the system processes a batch of tokens $\mathcal{I}_{t}$, which may consist of tokens from different prompts, but cannot include multiple tokens from the same prompt, as token generation is sequential. To ensure feasibility, the total memory usage in any batch must not exceed $M$, i.e.,
$\sum_{i\in\mathcal{I}_{t}}(s_{i}+a_{i})\leq M,$
where $a_{i}$ denotes the index of the token from request $i$ included in the batch.
Our objective is to minimize the total end-to-end latency, defined as the sum of completion times for all requests. Since all prompts arrive at $t=0$, the end-to-end latency of request $i$ is equal to the time at which its final token is processed. Letting $c_{i}(\Lambda)$ denote the completion time of request $i$ under schedule $\Lambda$, the total end-to-end latency is
$\text{TEL}(\Lambda;\mathcal{I})=\sum_{i\in[n]}c_{i}(\Lambda).$
To compute the optimal schedule, Jaillet et al. (2025) proposes an integer linear program (ILP). However, the ILP is computationally intractable in real-time inference settings, where scheduling decisions must be made within milliseconds.
Thus, our goal is to design efficient approximation algorithms. Let Optimal denote the optimal schedule with the smallest end-to-end latency for request set $\mathcal{I}$. For any scheduling algorithm $\Lambda$, we evaluate its performance using the competitive ratio:
$\text{CR}(\Lambda)=\sup_{\mathcal{I}}\frac{\text{TEL}(\Lambda;\mathcal{I})}{\text{TEL}(\text{{Optimal}};\mathcal{I})}.$
The competitive ratio measures the worst-case performance of the algorithm relative to the optimum. It is always at least 1, with smaller values indicating better performance. Our objective is to develop algorithms with low competitive ratios that are practical for real-time deployment.
2.1 Scheduling Challenges with Long Prompts
In Jaillet et al. (2025), the authors propose a memory-constrained shortest-first scheduling algorithm (MC-SF). The algorithm prioritizes requests with smaller output lengths $o_{i}$. When considering whether to add a prompt to the current batch, the algorithm precomputes the memory usage at each future time step to determine whether the addition would violate the memory constraint at any point in time. Formally, suppose a subset $U\subset R_{t}$ is considered for inclusion in the batch at time $t$. Define $t_{\max}(U):=\max_{i\in U}\{t+o_{i}\}$ as the latest possible completion time for jobs in $U$, assuming processing begins at time $t$. To ensure that the memory constraint is satisfied throughout the interval $t,t_{\max}(U)]$, the algorithm verifies that the total memory usage at each intermediate time $t^{\prime}\in[t,t_{\max}(U)]$ remains within the KV cache capacity $M$. Specifically, the memory-constrained shortest-first algorithm checks the following condition to determine whether including the batch $U$ is feasible at time $t$:
$\sum_{i\in S^{(t)}}(s+t^{\prime}-p_{i})\cdot\mathbbm{1}_{{o_{i}\geq t^{\prime}-p_{i}}}+\sum_{i\in U}(s+t^{\prime}-t)\cdot\mathbbm{1}_{{o_{i}\geq t^{\prime}-t}}\leq M,\quad\forall t^{\prime}\in[t,t_{\max}(U)].$ (1)
Here, $S^{(t)}$ denotes the set of jobs already in progress at time $t$, and $p_{i}$ represents the most recent start time of job $i$. The first summation accounts for the memory consumption of ongoing jobs that may still be active at time $t^{\prime}$, while the second summation captures the memory usage of new jobs in $U$ starting at time $t$. The indicator functions ensure that only those tokens still being generated at time $t^{\prime}$ contribute to the total memory usage.
Jaillet et al. (2025) demonstrates the strong theoretical and empirical performance of Algorithm MC-SF under the simplifying assumption that $s_{i}=s$ for all $i\in[n]$. This assumption makes the Algorithm MC-SF particularly natural: since $s_{i}$ determines the base memory usage, $o_{i}$ governs the processing time, and $s_{i}+o_{i}$ defines the peak memory usage, requests with smaller $o_{i}$ require less peak memory and complete faster. Thus, under uniform $s_{i}$, prioritizing shorter jobs both improves throughput and better utilizes memory.
However, the assumption of constant $s_{i}$ may not hold in practice, especially when the system must handle a mix of long and short prompts, leading to significant variability in input sizes. Relaxing this assumption introduces substantial complexity. For instance, consider two requests $i$ and $j$ where $s_{i}<s_{j}$ but $o_{i}>o_{j}$: request $i$ consumes less memory per time unit but runs longer, while request $j$ uses more memory but finishes quickly. In such cases, it is unclear which request should be prioritized, as the trade-off between memory consumption and completion time becomes nontrivial. Next, we formally show that when $s_{i}$ is allowed to vary, the competitive ratio of Algorithm MC-SF becomes unbounded.
Theorem 2.1
Suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then Algorithm MC-SF achieves an unbounded competitive ratio. Specifically, as $M\to\infty$, we have
$\mathrm{CR}(\textup{{MC-SF}})\to\infty.$
The proof of Theorem 2.1 can be found in Appendix 8. Theorem 2.1 illustrates that relaxing the assumption $s_{i}=s$ for all $i\in[n]$ can lead to many instances where existing scheduling algorithms perform poorly. In Appendix 8, we also show that if we do shortest-first according to $s_{i}+o_{i}$, the competitive ratio is still unbounded. Furthermore, the following theorem establishes that, without the assumption $s_{i}=s$, the problem of minimizing total end-to-end latency is NP-hard [^1] It remains an open question that either minimizing total end-to-end latency is NP-hard or not under the assumption $s_{i}=s$..
Theorem 2.2
Under the model suggested in Jaillet et al. (2025), suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then minimizing the total end-to-end latency is NP-hard.
We prove Theorem 2.2 by reduce the problem to a 3-Partition problem, and the rigorous proof can be found in Appendix 8. For readers interested in alternative objectives, we include in Appendix 8 a result (Theorem 8.5) showing that, under the same model, the problem remains NP-hard when the objective is to minimize the makespan. While this result is not central to the main focus of the paper, it further highlights the computational challenges of scheduling in this setting.
3 Algorithm Description
This section presents our novel scheduling algorithm designed to minimize total end-to-end latency. The algorithm dynamically schedules requests by balancing memory constraints with a quality metric that optimizes both batch concurrency and response length efficiency.
At the core of our approach is the quality metric $F(\mathcal{X})$ for any request set $\mathcal{X}$, defined as
$F(\mathcal{X})=\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|^{2}},$
where $\mathcal{X}$ is a set of requests, and $|\mathcal{X}|$ is the cardinality of the request set.
Smaller values of $F(\mathcal{X})$ indicate higher scheduling priority. This metric fundamentally improves upon shortest-first methods by naturally balancing batch size and response lengths. To illustrate its operational intuition, we begin with a concrete numerical example that demonstrates how Sorted-F outperforms shortest-first scheduling (MC-SF):
Example 3.1
Consider a scenario with KV cache memory $M=64$ and two request types:
Table 1: Request Types for Memory-Constrained Scheduling
Type Prompt Size Response Length Requests Number
1 63 1 1
2 1 2 21
Under MC-SF (which prioritizes shorter outputs), Type 1 would be scheduled first since $o_{1}=1<o_{2}=2$. After Type 1 completes at time $t=1$, Type 2 requests are processed in batches. With memory constraint $M=64$, each batch can accommodate $\lfloor 64/(1+2)\rfloor=21$ Type 2 requests simultaneously. The total end-to-end latency (TEL) is
$\text{TEL}(\text{{MC-SF}}\,;\mathcal{I})=\underbrace{1}_{\text{Type 1}}+\underbrace{21\times 3}_{\text{Type 2}}=64.$
In contrast, Sorted-F computes the F-metric for each type:
$F(\text{Type 1})$ $=\frac{1}{1^{2}}=1$
$F(\text{Type 2})$ $=\frac{2\times 21}{21^{2}}=\frac{42}{441}\approx 0.095.$
Since F(Type 2)<F(Type 1)F($\text${Type 2})<F($\text${Type 1}), Sorted-F prioritizes Type 2 first. Type 2 completes at $t=2$, after which Type 1 is processed at $t=3$. The TEL is
$\text{TEL}(\text{{Sorted-F}}\,;\mathcal{I})=\underbrace{21\times 2}_{\text{Type 2}}+\underbrace{3}_{\text{Type 1}}=45.$
This represents a $29.7\%$ reduction in TEL compared to MC-SF, demonstrating how Sorted-F’s F-metric effectively balances batch size and processing time to minimize latency.
Building on this concrete demonstration, we now generalize the intuition through a symbolic example that shows why the F-metric naturally leads to optimal scheduling decisions:
Example 3.2
Consider a simple scenario with two batches of requests where requests are homogeneous within each batch but heterogeneous across batches, as shown in Table 2.
Table 2: Batch Characteristics and Quality Metrics
Batch Prompt Size Response Length Requests Number Quality Metric
1 $s_{1}$ $o_{1}$ $n_{1}$ $o_{1}/n_{1}$
2 $s_{2}$ $o_{2}$ $n_{2}$ $o_{2}/n_{2}$
Assume that both batches have the same total memory requirement: $n_{1}(s_{1}+o_{1})=n_{2}(s_{2}+o_{2})=M$. The scheduler must decide which batch to process first.
If batch 1 is prioritized, it completes at time $o_{1}$, contributing $o_{1}n_{1}$ to the total end-to-end latency (TEL). Batch 2 then completes at time $o_{1}+o_{2}$, contributing $o_{2}n_{2}+o_{1}n_{2}$ to TEL. Thus, the TEL is $o_{1}n_{1}+o_{2}n_{2}+o_{1}n_{2}$.
Conversely, if batch 2 is prioritized, it completes at time $o_{2}$, contributing $o_{2}n_{2}$ to TEL. Batch 1 then completes at time $o_{1}+o_{2}$, contributing $o_{1}n_{1}+o_{2}n_{1}$ to TEL. The TEL in this case is $o_{1}n_{1}+o_{2}n_{2}+o_{2}n_{1}$.
The difference between the two schedules lies in the cross terms: $o_{1}n_{2}$ versus $o_{2}n_{1}$. Therefore, if $o_{1}/n_{1}<o_{2}/n_{2}$, then $o_{1}n_{2}<o_{2}n_{1}$, so prioritizing batch 1 minimizes TEL. Otherwise, prioritizing batch 2 minimizes TEL. Crucially, this optimal decision rule precisely aligns with selecting the batch with smaller $F(\mathcal{X})$, since $F(\mathcal{X})=o_{i}/n_{i}$ for each batch $i$. This demonstrates that the F-metric naturally guides the scheduler to the optimal choice.
Examples 3.1 and 3.2 demonstrate that the F-metric effectively guides scheduling decisions under different workloads. Beyond these illustrative cases, the metric exhibits valuable structural properties that contribute to robust performance in complex scenarios with varying input sizes ($s_{i}$) and output lengths ($o_{i}$).
A key challenge arises from this variability in output sizes: when requests are batched together, they do not necessarily finish processing at the same time. Under our model, once a request completes, its memory resources are freed, allowing new requests to be dynamically added to the batch. However, if output sizes in a batch vary too much, requests finish asynchronously, forcing the scheduler to switch from batch selection to incremental request insertion.
To address this, one such property—formalized in Lemma 3.3—ensures that within any selected batch, no single request’s output length dominates the average. This balancing characteristic prevents straggler effects and helps maintain predictable processing times, while also providing the analytical leverage needed for our competitive ratio guarantee in Section 4.
Lemma 3.3
For any batch $\mathcal{X}$ selected by the algorithm, let $o_{\max}=\max_{r_{i}\in\mathcal{X}}o_{i}$ and $\bar{o}=\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|}$. Then,
$\bar{o}>\frac{1}{2}\cdot o_{\max}.$
Proof 3.4
Proof of Lemma 3.3: From the algorithm’s selection criterion, any batch $\mathcal{X}$ must satisfy
$\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}-o_{\max}}{(|\mathcal{X}|-1)^{2}}\geq\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|^{2}}.$
Rearranging the terms yields
$\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|}\geq\frac{|\mathcal{X}|}{2|\mathcal{X}|-1}\cdot o_{\max}>\frac{1}{2}\cdot o_{\max}.$
[…]
Lemma 3.3 ensures that no single request’s output length dominates the batch average, making processing times predictable even under significant workload heterogeneity.
After having introduced the intuition and property behind the F-metric, we start to describe our scheduling algorithm, Sorted-F. The algorithm itself operates in two sequential phases, executed over discrete timesteps $t$. It tracks three request sets: pending requests $\mathcal{R}^{(t)}$, active requests $\mathcal{V}^{(t)}$, and newly initiated requests $\mathcal{U}^{(t)}$. The memory consumption at any future timestep $t^{*}$ is calculated using
$M(\mathcal{X},t^{*})=\sum_{r_{i}\in\mathcal{X}}\left(s_{i}+t^{*}-p_{i}\right)\cdot\mathds{1}_{\{o_{i}\geq t^{*}-p_{i}\}},$
which accounts for the cumulative growth of KV cache entries as tokens are generated. The complete procedure is formalized in Algorithm 1.
Algorithm 1 Sorted-F
1:Phase 1: Batch Construction
2:[$\mathcal{I}^{\prime}\leftarrow$[]
3:while $\mathcal{I}$ do
4:  $\mathcal{X}^{*}\leftarrow\arg\min_{\mathcal{X}\subseteq\mathcal{I}}(F(\mathcal{X}),-|\mathcal{X}|)$ subject to $M(\mathcal{X},p_{i}+o_{i})\leq M,\forall r_{i}\in\mathcal{X}$
5:  Sort $r_{i}\in\mathcal{X}^{*}$ in ascending order of $o_{i}$
6:  $\mathcal{I}^{\prime}\leftarrow\mathcal{I}^{\prime}+\mathcal{X}$, $\mathcal{I}\leftarrow\mathcal{I}-\mathcal{X}$
7:end while
8:Phase 2: Scheduling Execution
[…]
10:while $\mathcal{R}^{(t)}+\mathcal{V}^{(t)}$ do
11:  $t\leftarrow t+1$, $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}-\mathcal{U}^{(t-1)}$, $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t-1)}+\mathcal{U}^{(t-1)}$, $\mathcal{U}^{(t)}\leftarrow[]$
12:  for $r_{i}\in\mathcal{R}^{(t)}$ in sorted order do
13:   if $M(\mathcal{V}^{(t)}+\mathcal{U}^{(t)},p_{j}+o_{j})\leq M,\forall r_{j}\in\mathcal{V}^{(t)}+\mathcal{U}^{(t)}+[r_{i}]$ then
14:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}+[i]$
15:   else
16:     Break
17:   end if
18:  end for
[…]
20:end while
In Phase 1 (line 1-7), Sorted-F iteratively constructs request batches $\mathcal{X}^{*}$ that minimize $F(\mathcal{X})$ while respecting memory constraints $M(\mathcal{X},p_{i}+o_{i})\leq M$. Selected requests in a same batch are sorted by ascending $o_{i}$. The sorting criterion ensures strict ordering between batches: the request with maximal output length $o_{i}^{\max}$ in batch $B_{k}$ immediately precedes the request with minimal output length $o_{i}^{\min}$ in batch $B_{k+1}$. This generates a sorted request set $\mathcal{I}^{\prime}$.
Phase 2 (line 8-20) handles real-time token generation by: dynamically updating request sets at each timestep; admitting new requests in $o_{i}$-sorted order only if they preserve memory constraints at all future timesteps; and processing one token per active request concurrently.
The algorithm’s design provides significant advantages. First, the output length bound $o_{i}\leq 2\cdot\bar{o}$ ensures predictable batch processing times. Second, the F-metric prevents both small-$s_{i}$ large-$o_{i}$ requests from monopolizing early resources and large-$s_{i}$ small-$o_{i}$ requests from creating fragmentation. Third, the two-phase structure enables efficient hardware implementation by decoupling optimization from execution. Through these mechanisms, our approach effectively addresses the scheduling pathology from first-come-first-serve or shortest-first demonstrated in Theorem 2.1.
4 Proof of Constant Competitive Ratio
In this section, we prove the following theorem:
Theorem 4.1
Sorted-F achieves a constant competitive ratio upper bounded by $48$ against the optimal schedule Optimal, i.e.,
$\mathrm{CR}(\textup{{Sorted-F}})<48.$
To prove Theorem 4.1, we transform Sorted-F sequentially: $\textup{{Sorted-F}}\to\textup{{Sorted-F}}_{\mathrm{separate}}$ (Theorem 4.2) $\to\textup{{Sorted-F}}_{\mathrm{group}}$ (Theorem 4.5) $\to\textup{{Sorted-F}}_{\mathrm{align}}$ (Theorem 4.8). At each step, we deliberately modify the schedule’s structure, but crucially, we prove that the cumulative latency increase remains within constant factors. In parallel, we restructure Optimal into $\textup{{Optimal}}_{\mathrm{align}}$ (Theorem 4.10). The proof culminates in a direct comparison between $\textup{{Sorted-F}}_{\mathrm{align}}$ and $\textup{{Optimal}}_{\mathrm{align}}$ (Theorem 4.12).
Before delving into the technical steps, we clarify the asymptotic regime in which the analysis is carried out. Throughout the proof, we consider a large-instance setting where the request set $\mathcal{I}$ is sufficiently large relative to the memory budget $M$. Recall from Section 2 that each request has size $s_{i},o_{i}=\epsilon(M)$, meaning that individual requests consume a vanishing fraction of memory. W.L.O.G, we further assume that the total number of requests $n=n(M)$ satisfies $\frac{n(M)\,\epsilon(M)}{M}\;\longrightarrow\;\infty$ as $M\to\infty$. Otherwise, by processing the finite amount of requests, any scheduling algorithm can have a constant competitive ratio. In words, the request set is effectively infinite, and its cardinality grows much faster than the maximum number of requests that can fit into a single batch of capacity $M$. This ensures that the contribution of any fixed group becomes negligible compared with the “tail” of remaining requests, thereby justifying the asymptotic bounds used in later arguments.
With this standing assumption in place, we now begin the first step of the proof. Let $\{\mathcal{X}_{k}\}$ denote the sequence of batches iteratively generated in Phase 1 of Sorted-F. To construct $\textup{{Sorted-F}}_{\mathrm{separate}}$, we process $\{\mathcal{X}_{k}\}$ one at a time. Specifically, for any batch $\mathcal{X}_{k}$, all requests in it start simultaneously, and the subsequent batch $\mathcal{X}_{k+1}$ does not begin until all requests in $\mathcal{X}_{k}$ completed. This sequential batch processing simulates a worst-case scenario by ensuring that no request from subsequent batches is initiated until the preceding batch has fully finished.
Let the requests in $\mathcal{X}_{k}$, ordered by increasing response length, be denoted as $r_{k,1},r_{k,2},\ldots,r_{k,n_{k}}$, where $n_{k}=|\mathcal{X}_{k}|$. Then the process of transforming Sorted-F into $\textup{{Sorted-F}}_{\mathrm{separate}}$ is depicted in Figure 1.
Figure 1: From Sorted-F to $\textup{{Sorted-F}}_{\mathrm{separate}}$
Theorem 4.2 characterizes the relationship between the total end-to-end latency of Sorted-F and that of $\textup{{Sorted-F}}_{\mathrm{separate}}$.
Theorem 4.2
For any $\mathcal{I}$, the total end-to-end latency of Sorted-F is bounded above by that of $\text{{Sorted-F}}_{\text{separate}}$, i.e.,
$\mathrm{TEL}(\textup{{Sorted-F}};\mathcal{I})\leq\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{separate}};\mathcal{I}).$
Proof 4.3
Proof of Theorem 4.2: Let $p_{k,j}$ and $p^{\prime}_{k,j}$ denote the initiation times of request $r_{k,j}$ under Sorted-F and $\textup{{Sorted-F}}_{\mathrm{separate}}$, respectively. To prove the theorem, we show that for all batches $\mathcal{X}_{k}$ and all $j\in\{1,\dots,n_{k}\}$,
$p_{k,j}\leq p^{\prime}_{k,j}.$
This inequality holds because $\textup{{Sorted-F}}_{\mathrm{separate}}$ enforces sequential batch processing: no request in $\mathcal{X}_{k+1}$ starts until all requests in $\mathcal{X}_{k}$ complete. Thus, $\textup{{Sorted-F}}_{\mathrm{separate}}$ delays requests relative to Sorted-F, which allows overlapping batch execution.
To formalize this, we analyze the memory usage constraints. Consider modifying $\textup{{Sorted-F}}_{\mathrm{separate}}$ by iteratively advancing the initiation times $p^{\prime}_{k,j}$ to their earliest feasible points, while respecting memory limits. For each $\mathcal{X}_{k}$, process requests in order $j=1,2,\cdots,n_{k}$. After advancing $r_{k,j-1}$ ($j\geq 2$), the maximum memory usage in the interval $\max\{p^{\prime}_{k,1},p_{k,j-1}+o_{k,j-1}\},p^{\prime}_{k,j}+o_{k,j}]$ is
$\sum_{i=j}^{n_{k}}s_{i}+o_{k,n_{k}}\cdot(n_{k}-j+1).$
Advancing $r_{k,j}$ reduces this usage to
$\sum_{i=j}^{n_{k}}s_{i}+(o_{k,n_{k}}-(p^{\prime}_{k,j}-p_{k,j}))^{+}\cdot(n_{k}-j+1)\leq\sum_{i=j}^{n_{k}}s_{i}+o_{k,n_{k}}\cdot(n_{k}-j+1).$
[…]
Definition 4.4
A batch $\mathcal{X}_{k}$ nearly saturates memory $M$ if
$\sum_{r_{i}\in\mathcal{X}_{k}}(s_{i}+o_{i})>M-\epsilon(M),$
where $\epsilon(M)$ is a negligible term asymptotically smaller than $M$.
Given Definition 4.4, we present Theorem 4.5 to establish an upper bound of the total end-to-end latency of $\textup{{Sorted-F}}_{\mathrm{separate}}$.
Theorem 4.5
For any $\mathcal{I}$, there exists a grouping strategy $\text{{Sorted-F}}_{\text{group}}$ that aggregates $\{\mathcal{X}_{k}\}$ into larger groups $\{\mathcal{Y}_{m}\}$ such that the terminal batch of each group nearly saturates memory $M$. Additionally, the total end-to-end latency of $\text{{Sorted-F}}_{\text{separate}}$ is bounded above by
$\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{separate}};\mathcal{I})<4\cdot\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{group}};\mathcal{I}).$
Proof 4.6
Proof of Theorem 4.5: We construct aggregated groups $\{\mathcal{Y}_{m}\}$ via
$\mathcal{Y}_{m}=\mathcal{X}_{b_{m-1}+1}+\mathcal{X}_{b_{m-1}+2}+\dots+\mathcal{X}_{b_{m}},$
where $b_{m}=\sum_{j=1}^{m}a_{j}$ tracks cumulative batches, $a_{j}$ counts batches in $\mathcal{Y}_{j}$, and the addition represents concurrent initiation of all requests in constituent batches. Figure 2 illustrates this transformation from $\textup{{Sorted-F}}_{\mathrm{separate}}$ to $\textup{{Sorted-F}}_{\mathrm{group}}$.
Figure 2: From $\textup{{Sorted-F}}_{\mathrm{separate}}$ to $\textup{{Sorted-F}}_{\mathrm{group}}$
Each group $\mathcal{Y}_{m}$ is constructed such that $b_{m}$ is the smallest integer $k>b_{m-1}$ satisfying the inequality
$\frac{\sum_{j=1}^{n_{k}}o_{k,j}+o_{k+1,1}}{(n_{k}+1)^{2}}\leq\frac{\sum_{j=1}^{n_{k}}o_{k,j}}{n_{k}^{2}},$
which implies that including $r_{b_{m}+1,1}$ in $\mathcal{X}_{b_{m}}$ would violate the memory constraint $M$. If this were not the case, the request $r_{b_{m}+1,1}$ should have been incorporated into $\mathcal{X}_{b_{m}}$ to either reduce the objective value $F(\mathcal{X}_{b_{m}})$ or increase the batch size $|\mathcal{X}_{b_{m}}|$ while preserving $F(\mathcal{X}_{b_{m}})$. However, such inclusion would contradict the optimal selection criterion of Sorted-F. We therefore conclude that $\mathcal{X}_{b_{m}}$ must nearly saturate the memory $M$ according to Definition 4.4.
Next, we present Lemma 4.7 to characterize the geometric relationship between processing times of batches within each group, where the processing time of a batch is defined as the longest response length among all its requests.
Lemma 4.7
For any group $\mathcal{Y}_{m}$ and any batch index $i\in\{1,\dots,a_{m}\}$, the longest processing time $o_{b_{m-1}+i,n_{b_{m-1}+i}}$ of batch $\mathcal{X}_{b_{m-1}+i}$ satisfies
$o_{b_{m-1}+i,n_{b_{m-1}+i}}\leq\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-i}{2}\rfloor}\cdot o_{b_{m},n_{b_{m}}},$
where $o_{b_{m},n_{b_{m}}}$ is the longest processing time in the terminal batch $\mathcal{X}_{b_{m}}$ of group $\mathcal{Y}_{m}$.
The proof of Lemma 4.7 appears in Appendix 9. By Lemma 4.7, we can derive
$\sum_{i=1}^{a_{m}}o_{b_{m-1}+i,n_{b_{m-1}+i}}$ $\leq\sum_{i=1}^{a_{m}}\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-i}{2}\rfloor}\cdot o_{b_{m},n_{b_{m}}}$
$\leq 2\cdot\frac{1-\left(\frac{1}{2}\right)^{\lfloor\frac{a_{m}-1}{2}\rfloor}}{1-\frac{1}{2}}\cdot o_{b_{m},n_{b_{m}}}$
$\leq 4\cdot o_{b_{m},n_{b_{m}}}.$
Therefore, given any $\mathcal{I}$,
$\text{TEL}(\text{{Sorted-F}}_{\text{group}};\mathcal{I})$ $=\sum_{m=1}^{\infty}\left(o_{b_{m},n_{b_{m}}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$\geq\sum_{m=1}^{\infty}\left(\frac{1}{4}\cdot\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$=\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\left(\sum_{k=b_{m}+1}^{\infty}n_{k}+\epsilon\left(\sum_{k=b_{m}+1}^{\infty}n_{k}\right)\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$=\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\left(\sum_{k=b_{m}+1}^{\infty}n_{k}+\sum_{k=b_{m-1}+1}^{m}n_{k}\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$>\frac{1}{4}\cdot\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}+\sum_{l=b_{m-1}+1}^{b_{m}-1}\left(o_{l,n_{l}}\cdot\left(b_{m}-l\right)\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$>\frac{1}{4}\cdot\left(\sum_{m=1}^{\infty}\left(\sum_{l=b_{m-1}+1}^{b_{m}}o_{l,n_{l}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}+\sum_{l=b_{m-1}+1}^{b_{m}-1}\left(o_{l,n_{l}}\cdot\left(b_{m}-l\right)\right)\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}\right)$
$=\frac{1}{4}\cdot\text{TEL}(\text{{Sorted-F}}_{\text{separate}};\mathcal{I}).$
[…]
Next, we focus on the terminal batches $\{\mathcal{X}_{b_{m}}\}$. To construct $\text{{Sorted-F}}_{\text{align}}$, we replace each request’s original response length in $\mathcal{X}_{b_{m}}$ with the batch’s average response length:
$\bar{o}_{b_{m}}=\frac{\sum_{i=1}^{n_{b_{m}}}o_{b_{m},i}}{n_{b_{m}}}.$
This conversion from $\textup{{Sorted-F}}_{\mathrm{group}}$ to $\textup{{Sorted-F}}_{\mathrm{align}}$ is shown in Figure 3.
Figure 3: From $\textup{{Sorted-F}}_{\mathrm{group}}$ to $\textup{{Sorted-F}}_{\mathrm{align}}$
Theorem 4.8 compares the total end-to-end latency of $\text{{Sorted-F}}_{\text{align}}$ with $\text{{Sorted-F}}_{\text{group}}$.
Theorem 4.8
For any $\mathcal{I}$, the total end-to-end latency of $\text{{Sorted-F}}_{\text{group}}$ is bounded above by
$\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{group}};\mathcal{I})<2\cdot\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{align}};\mathcal{I}).$
Proof 4.9
Proof of Theorem 4.8: By Lemma 3.3,
$\bar{o}_{b_{m}}>\frac{1}{2}\cdot o_{b_{m},n_{b_{m}}}.$
Subsequently, given any $\mathcal{I}$,
$\text{TEL}(\text{{Sorted-F}}_{\text{align}};\mathcal{I})$ $=\sum_{m=1}^{\infty}\left(\bar{o}_{b_{m}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$>\sum_{m=1}^{\infty}\left(\frac{1}{2}\cdot o_{b_{m},n_{b_{m}}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}$
$\geq\frac{1}{2}\cdot\left(\sum_{m=1}^{\infty}\left(o_{b_{m},n_{b_{m}}}\cdot\sum_{k=b_{m}+1}^{\infty}n_{k}\right)+\sum_{k=1}^{\infty}\sum_{i=1}^{n_{k}}o_{k,i}\right)$
$=\frac{1}{2}\cdot\text{TEL}(\text{{Sorted-F}}_{\text{group}};\mathcal{I}).$
[…]
By definition 4.4, we derive a lower bound on the average memory utilization efficiency $\eta_{m}$ for any group $\mathcal{Y}_{m}$:
$\eta_{m}$ $\geq\frac{\sum_{r_{i}\in\mathcal{Y}_{m}}(2s_{i}+o_{i})}{2\cdot M}>\frac{M-\epsilon(M)}{2\cdot M}=\frac{1}{2}.$ (2)
Now, we introduce Theorem 4.10 to construct a lower bound of the total end-to-end latency of Optimal.
Theorem 4.10
For any $\mathcal{I}$, there exists a transformed schedule $\text{{Optimal}}_{\text{align}}$ with time-window scaling and response-length aligning that satisfies
$\mathrm{TEL}(\textup{{Optimal}}_{\mathrm{align}};\mathcal{I})\leq 6\cdot\mathrm{TEL}(\textup{{Optimal}};\mathcal{I}).$
Proof 4.11
Proof of Theorem 4.10: We construct $\text{{Optimal}}_{\text{align}}$ through an iterative grouping process consisting of two phases: first partitioning requests into temporal groups $\{\mathcal{Z}_{g}\}$, then further decomposing each group into execution batches $\{\mathcal{W}_{h}\}$. This process of transforming Optimal into $\textup{{Optimal}}_{\mathrm{align}}$ is demonstrated in Figure 4.
Figure 4: From Optimal to $\textup{{Optimal}}_{\mathrm{align}}$
The construction begins with the initial group $\mathcal{Z}_{1}$, defined as follows:
Let $\mathcal{A}_{1}$ denote the set of all requests initiated at time 0. We define
$\delta_{1}=\max\{p_{i}+o_{i}\mid r_{i}\in\mathcal{A}_{1}\},$
which establishes the maximum processing duration among initial requests. This yields our first group:
$\mathcal{Z}_{1}=\{r_{i}\in\mathcal{I}\mid p_{i}+o_{i}\leq\delta_{1}\},$
containing all requests whose total duration falls within this initial temporal bound.
For subsequent groups ($g\geq 2$), we define $\delta_{g}$ recursively as
$\delta_{g}=\max\{p_{i}+o_{i}\mid r_{i}\in\mathcal{A}_{g}\}-\sum_{j=1}^{g-1}\delta_{j},$
where $\mathcal{A}_{g}$ consists of requests initiated but not yet completed by time $\sum_{j=1}^{g-1}\delta_{j}$. The $g$-th group $\mathcal{Z}_{g}$ is constructed as
$\mathcal{Z}_{g}=\bigg{\{}r_{i}\in\mathcal{I}\,\bigg{|}\,\sum_{j=1}^{g-1}\delta_{j}<p_{i}+o_{i}\leq\sum_{j=1}^{g}\delta_{j}\bigg{\}}.$
The transformation begins by delaying the execution of each group $\mathcal{Z}_{g}$ ($g\geq 2$) by $5\sum_{j=2}^{g-1}\delta_{j}$ time units while maintaining the internal timing relationships among all requests within the group. For $g\geq 2$, we define the adjusted cutting time as
$t_{g}=6\sum_{j=2}^{g-1}\delta_{j},$
effectively scaling the original temporal partitioning by a factor of 6.
Next, we decompose each group $\mathcal{Z}_{g}$ into smaller sub-batches $\{\mathcal{W}_{h}\}$ through the following partition:
$\mathcal{Z}_{g}=\mathcal{W}_{d_{g-1}+1}+\mathcal{W}_{d_{g-1}+2}+\cdots+\mathcal{W}_{d_{g}},$
where $d_{g}=\sum_{j=1}^{g}c_{j}$ counts the cumulative batches up to group $g$, with $c_{j}$ denoting the number of batches in $\mathcal{Z}_{j}$. The decomposition proceeds by first sorting all requests in $\mathcal{Z}_{g}$ by their original initiation times. Tnhen, for each subsequent batch $\mathcal{W}_{d_{g-1}+k}$, we select the maximal subset of remaining requests with the earliest initiation times while ensuring the total memory requirement does not exceed capacity $M$. This process repeats until all requests in $\mathcal{Z}_{g}$ are assigned to batches. This process guarantees that all batches except possibly the final one ($\mathcal{W}_{d_{g}}$) nearly saturate memory $M$ by Definition 4.4.
We now replace each request’s original response length in $\mathcal{W}_{h}$ with the batch’s average response length:
$\bar{o}_{h}=\frac{1}{n_{h}}\sum_{i=1}^{n_{h}}o_{h,i},$
where $n_{h}$ is the number of requests in batch $\mathcal{W}_{h}$. Let $\text{OPT}_{\text{transform}}$ denote the total end-to-end latency after this transformation.
By Definition 4.4, we derive a lower bound on the memory utilization efficiency $\eta_{h}$ for any non-final batch $\mathcal{W}_{h}$ (where $h\neq d_{g}$):
$\eta_{h}\geq\frac{\sum_{r_{i}\in\mathcal{W}_{h}}(2s_{i}+o_{i})}{2M}>\frac{M-\epsilon(M)}{2M}=\frac{1}{2}.$ (3)
Given that the original processing window for $\mathcal{Z}_{g}$ satisfies $\Delta t_{g}<\delta_{g-1}+\delta_{g}$, inequality 3 implies that the main batches $\mathcal{W}_{d_{g-1}+1},\ldots,\mathcal{W}_{d_{g}-1}$ complete within $2(\delta_{g-1}+\delta_{g})$ time units. Furthermore, the final batch $\mathcal{W}_{d_{g}}$, with all response lengths bounded by $\delta_{g-1}+\delta_{g}$, completes within $\delta_{g-1}+\delta_{g}$ time units. Consequently, the complete batch sequence $\{\mathcal{W}_{d_{g-1}+k}\}_{k=1}^{c_{g}}$ is guaranteed to finish within the time interval $t_{g}-3\delta_{g-1},t_{g}+3\delta_{g}]$.
Therefore, given any $\mathcal{I}$, we establish a lower bound on Optimal’s total end-to-end latency through the following derivation:
$\text{TEL}(\text{{Optimal}}_{\text{transform}};\mathcal{I})$ $=\sum_{h=1}^{\infty}\left(\bar{o}_{h}\cdot\sum_{j=h+1}^{\infty}n_{j}\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$=\sum_{g=1}^{\infty}\sum_{h=d_{g-1}+1}^{d_{g}}\left(\bar{o}_{h}\cdot\left(\sum_{j=h+1}^{d_{g}}n_{j}+\sum_{j=d_{g}+1}^{\infty}n_{j}\right)\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$=\sum_{g=1}^{\infty}\sum_{h=d_{g-1}+1}^{d_{g}}\left(\bar{o}_{h}\cdot\left(\epsilon\left(\sum_{j=d_{g}+1}^{\infty}n_{j}\right)+\sum_{j=d_{g}+1}^{\infty}n_{j}\right)\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$=\sum_{g=1}^{\infty}\left(\sum_{h=d_{g-1}+1}^{d_{g}}\bar{o}_{h}\cdot\sum_{j=d_{g}+1}^{\infty}n_{j}\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$\leq\sum_{g=1}^{\infty}\left(3\cdot(\delta_{g-1}+\delta_{g})\cdot\sum_{j=d_{g}+1}^{\infty}n_{j}\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$=3\cdot\sum_{g=1}^{\infty}\left(\delta_{g}\cdot\left(\sum_{j=d_{g}+1}^{\infty}n_{j}+\sum_{j=d_{g+1}+1}^{\infty}n_{j}\right)\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$\leq 6\cdot\sum_{g=1}^{\infty}\left(\delta_{g}\cdot\sum_{j=d_{g}+1}^{\infty}n_{j}\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}$
$\leq 6\cdot\left(\sum_{g=1}^{\infty}\left(\delta_{g}\cdot\sum_{j=d_{g}+1}^{\infty}n_{j}+\sum_{r_{l}\in\mathcal{Z}_{g}}(p_{i}-t_{g})\right)+\sum_{h=1}^{\infty}\sum_{i=1}^{n_{h}}o_{h,i}\right)$
$=6\cdot\text{TEL}(\text{{Optimal}};\mathcal{I}).$
[…]
Finally, Theorem 4.12 provides the critical linkage between $\text{{Sorted-F}}_{\text{align}}$ and $\text{{Optimal}}_{\text{align}}$.
Theorem 4.12
For any $\mathcal{I}$, the total end-to-end latency of $\text{{Sorted-F}}_{\text{align}}$ is upper bounded by the total end-to-end latency of $\text{{Optimal}}_{\text{align}}$, i.e.,
$\mathrm{TEL}(\textup{{Sorted-F}}_{\mathrm{align}};\mathcal{I})\leq\mathrm{TEL}(\textup{{Optimal}}_{\mathrm{align}};\mathcal{I}).$
Proof 4.13
Proof of Theorem 4.12 The key insight is that minimizing total end-to-end latency for a fixed makespan requires maximizing throughput in the earliest possible time intervals. This relationship is captured precisely by our throughput metric $\Phi(\mathcal{X}_{k})=n_{k}/\bar{o}_{k}$, which measures requests completed per unit time in batch $\mathcal{X}_{k}$. Notably, this metric is exactly the reciprocal of our quality metric $F(\mathcal{X}_{k})$, establishing a direct connection between our optimization objective and scheduling efficiency.
The algorithm $\text{{Sorted-F}}_{\text{align}}$ is designed to explicitly optimize for early throughput maximization by minimizing $F(\mathcal{X}_{k})$. For each batch $\mathcal{X}_{k}$, given the fixed prior batches $\mathcal{X}_{1},\mathcal{X}_{2},...,\mathcal{X}_{k-1}$, $\text{{Sorted-F}}_{\text{align}}$ achieves the minimal possible $F(\mathcal{X}_{k})$, thereby maximizing $\Phi(\mathcal{X}_{k})$.
Furthermore, the memory utilization efficiency comparison reveals a significant advantage: in Algorithm $\text{{Sorted-F}}_{\text{align}}$, each group $\mathcal{Y}_{m}$ achieves memory utilization efficiency $\eta_{m}>1/2$ (Inequality 2), whereas in $\text{{Optimal}}_{\text{align}}$, the scaled allocation limits the efficiency to $\eta_{[t_{g},t_{g+1}]}\leq 1/6$. This efficiency gap directly translates to a strictly smaller makespan for $\text{{Sorted-F}}_{\text{align}}$ compared to $\text{{Optimal}}_{\text{transform}}$.
[…]
Building upon these results, we now establish Theorem 4.1.
Proof 4.14
Proof of Theorem 4.1: Given any $\mathcal{I}$, we combine the guarantees from Theorems 4.2, 4.5, 4.8, 4.10, and 4.12 to derive the following chain of inequalities for the total end-to-end latency:
$\text{TEL}(\text{{Sorted-F}};\mathcal{I})\leq\text{TEL}(\text{{Sorted-F}}_{\text{separate}};\mathcal{I})<4\cdot\text{TEL}(\text{{Sorted-F}}_{\text{group}};\mathcal{I})$
$<8\cdot\text{TEL}(\text{{Sorted-F}}_{\text{align}};\mathcal{I})\leq 8\cdot\text{TEL}(\text{{Optimal}}_{\text{align}};\mathcal{I})\leq 48\cdot\text{TEL}(\text{{Optimal}};\mathcal{I}),$
which implies that the competitive ratio of Sorted-F is bounded by
$\text{CR}(\text{{Sorted-F}})=\sup_{\mathcal{I}}\frac{\text{TEL}(\text{{Sorted-F}};\mathcal{I})}{\text{TEL}(\text{{Optimal}};\mathcal{I})}<48.$
[…]
5 Approximation Methods for Sorted-F
As described in Section 3, the Phase 1 scheduling of Sorted-F requires solving the combinatorial optimization problem $\mathcal{X}^{*}=\arg\min_{\mathcal{X}\subseteq\mathcal{I}}F(\mathcal{X})$ subject to memory constraints $M(\mathcal{X},p_{i}+o_{i})\leq M$. As exhaustive search is infeasible for large request sets, we develop four approximation methods spanning the scalability spectrum. Each method makes distinct trade-offs between optimality guarantees and computational complexity.
5.1 Exact Dynamic Programming
This algorithm employs dynamic programming to find the optimal request subset $\mathcal{X}^{*}\subseteq\mathcal{I}$ that minimizes the quality metric $F(\mathcal{X})$ within memory constraint $M$. The solution leverages a dual-component state representation:
State: $dp[k][m]$ $=\min\left\{\sum_{r_{i}\in\mathcal{X}}o_{i}:|\mathcal{X}|=k,\ \sum_{r_{j}\in\mathcal{X}}(s_{j}+o_{j})=m\right\}$
Path: $path[k][m]$ $=\mathcal{X}\text{ achieving }dp[k][m]$
Objective: $\mathcal{X}^{*}$ $=\arg\min_{\begin{subarray}{c}k\in[1,n]\\ m\leq M\end{subarray}}\left\{\frac{dp[k][m]}{k^{2}}\right\}$
The state $dp[k][m]$ captures the minimal sum of output lengths for any $k$-request batch consuming $m$ memory, while the path component $path[k][m]$ records the actual request indices achieving this value. This dual representation enables efficient exploration of the solution space and exact reconstruction of the optimal request set. The complete procedure is formalized in Algorithm 2, which can be found in Appendix 10.
The reverse traversal order is essential to prevent multiple inclusions of the same request while preserving the optimal substructure property. Path reconstruction occurs after all state transitions complete, converting stored indices into actual request objects. The $O(n^{2}M)$ complexity provides a practical exact solution for moderate-scale deployments where $n\leq 100$, establishing the fundamental benchmark for optimal scheduling decisions.
5.2 Scaled Dynamic Programming
This method provides a memory-efficient approximation to the exact combinatorial optimization problem. By quantizing memory usage values, it achieves polynomial complexity while maintaining $(1+\epsilon)$-optimality guarantees. The core innovation is a scaled representation of memory constraints:
$\hat{m}_{i}$ $=\left\lfloor\frac{s_{i}+o_{i}}{\lambda}\right\rfloor\quad\text{with}\quad\lambda=\frac{\epsilon M}{B}$
$\hat{M}$ $=\left\lfloor\frac{M}{\lambda}\right\rfloor,$
where $B$ controls precision. This reduces the state space dimension by factor $\lambda$, transforming the optimization to
$\mathcal{X}^{*}_{approx}=\arg\min_{\mathcal{X}\subseteq\mathcal{I}}$ $\frac{\sum_{r_{i}\in\mathcal{X}}o_{i}}{|\mathcal{X}|^{2}}$
subject to $\sum_{r_{i}\in\mathcal{X}}\hat{m}_{i}\leq\hat{M}$.
The complete pseudocode is provided in Algorithm 3 in Appendix 10.
The space-time complexity is $O(nB/\epsilon)$ versus $O(n^{2}M)$ for exact DP, enabling scaling to $n\leq 200$. The approximation quality exhibits a controllable trade-off: smaller $\epsilon$ values yield tighter bounds $F_{approx}\leq(1+\epsilon)F^{*}$ at the cost of $O(1/\epsilon)$ computation increase. This approach provides a theoretically grounded framework for balancing accuracy and efficiency in large-scale scenarios.
5.3 Local Swap Search
This heuristic algorithm efficiently refines an initial solution through iterative local improvements. Starting with a feasible batch $\mathcal{X}_{0}$ constructed by sorting requests in ascending order of $s_{i}+o_{i}$ and applying memory-constrained greedy selection, the method systematically explores request swaps to progressively enhance solution quality. The approach exploits the combinatorial structure of the optimization landscape while avoiding exhaustive search.
The core operation is pairwise exchange: for any $r_{out}\in\mathcal{X}$ and $r_{in}\notin\mathcal{X}$ satisfying
  1. 1.
    Memory feasibility: $\text{mem}+(s_{in}+o_{in})-(s_{out}+o_{out})\leq M$,
  2. 2.
    Quality improvement: $F(\mathcal{X}^{\prime})<F(\mathcal{X})$ with $\mathcal{X}^{\prime}=(\mathcal{X}\setminus\{r_{out}\})\cup\{r_{in}\}$.
Formally, the iterative refinement is described by
$\mathcal{X}_{k+1}=\begin{cases}\mathcal{X}_{k}&\nexists\text{ improving swap}\\[5.0pt] \mathcal{X}_{k}\oplus(r_{in},r_{out})&F(\mathcal{X}_{k}\oplus(r_{in},r_{out}))<F(\mathcal{X}_{k}),\end{cases}$
where $\oplus$ denotes the exchange operation. The algorithm terminates at local minima where no improving swaps exist. The complete procedure is formalized in Algorithm 4, which can be found in Appendix 10.
The $O(n^{2})$ complexity enables deployment at scale $n\leq 500$. While lacking global optimality guarantees, the method demonstrates strong practical performance, yielding high-quality solutions that approach optimum scheduling efficiency with significantly reduced computational overhead compared to exact methods.
5.4 Quantile Greedy Selection
This approach efficiently handles large-scale scheduling problems by combining statistical sampling with two-phase greedy selection. The method strategically limits outlier requests while optimizing memory utilization for maximum throughput:
Phase 1 (Core Selection): $\mathcal{X}_{1}$ $=\{r_{i}\in\mathcal{I}:s_{i}+o_{i}\leq Q_{p},\ o_{i}\leq Q_{o}\}$
Phase 2 (Capacity Filling): $\mathcal{X}$ $=\mathcal{X}_{1}\cup\left\{r_{j}\in\mathcal{I}\setminus\mathcal{X}_{1}:\text{mem}+s_{j}+o_{j}\leq M\right\},$
where quantile thresholds $Q_{p}$ and $Q_{o}$ are derived from workload sampling. The complete pseudocode is provided in Algorithm 5 in Appendix 10.
The $O(n)$ complexity ensures practical deployment for large-scale systems with $n\geq 1000$ requests. Phase 1 prioritizes predictable requests within distributional norms, while Phase 2 maximizes marginal utility of memory usage through the $o_{i}/(s_{i}+o_{i})$ efficiency metric. This combined approach maintains low computational overhead while effectively constraining outlier impact on scheduling performance.
5.5 Algorithm Selection Guidelines
The four approximation algorithms present distinct operational characteristics that enable optimized deployment across the scalability spectrum. System designers should consider the algorithmic trade-offs in terms of solution quality, computational efficiency, and applicability range when selecting schedulers for specific deployment scenarios. The comparative properties are formally summarized in Table 3, which serves as the primary reference for implementation decisions.
Table 3: Algorithmic Properties and Operational Ranges
Algorithm Optimality Complexity Recommended Scale
Exact Dynamic Programming Optimal $O(n^{2}M)$ $n\leq 100$
Scaled DP $(1+\epsilon)$-approximation $O(nB/\epsilon)$ $n\leq 200$
Local Swap Search Local optimum $O(n^{2})$ $n\leq 500$
Quantile Greedy Heuristic $O(n)$ $n\geq 1000$
Selection decisions should prioritize alignment between algorithm capabilities and deployment requirements, with accuracy-sensitive systems favoring exact methods despite computational costs, while latency-critical large-scale deployments benefit from linear-time heuristics. For moderate-scale systems with $200\leq n\leq 500$, the local search approach offers a balanced compromise, delivering near-optimal solutions with quadratic complexity that remains computationally manageable. Environmental factors including workload heterogeneity and available hardware resources further influence implementation choices; heterogeneous systems particularly benefit from the Quantile Greedy’s statistical sampling approach that constrains outlier impact through distribution-aware selection. When considering temporal performance profiles, designers should note that while worst-case complexity bounds determine theoretical scaling, practical implementations frequently achieve substantially better performance, particularly for the Local Swap method which empirically converges in constant iterations across diverse workloads. Hybrid implementations can dynamically switch algorithms based on real-time request volume to maintain optimal scheduling efficiency across varying load conditions. In the next section, we will compare the performance of these approximation methods on real-data numerical experiments.
6 Extensions and Discussions
Building on our earlier work—where we proposed the Sorted-F algorithm and established its constant-factor competitive ratio for this NP-hard problem—we now delve deeper into the problem’s underlying structure. Specifically, we analyze it through the lenses of integer programming (IP) and linear programming (LP), exploring alternative heuristics derived from these frameworks. This investigation not only broadens the theoretical understanding of the problem but also enables a systematic comparison between IP/LP-based approaches and our Sorted-F algorithm, shedding light on their relative strengths and practical trade-offs.
From Jaillet et al. (2025), the optimal schedule can be solved by the following integer programming formulation:
$\text{OPT-IP}=\min\,$ $\sum_{i\in[n]}\left(\sum_{t=\{0,\dots,\bar{T}\}}t\cdot x_{i,t}+o_{i}\right)$ (4)
s.t. $\sum_{t=\{0,\dots,\bar{T}\}}x_{i,t}=1,$ $\forall i\in[n]$ (5)
$\sum_{i=1}^{n}\sum_{k=\max\{0,t-o_{i}\}}^{t-1}(s_{i}+t-k)x_{i,k}\leq M,$ $\forall t\in[\bar{T}]$ (6)
$x_{i,t}\in\{0,1\},$ $\forall i\in[n],\forall t\in[\bar{T}]$ (7)
where $\bar{T}$ is the upper bound of the time horizon. The decision variables are $x_{i,t}$. If $x_{i,t}=1$, the request $i$ should start at time $t$. Equation (5) guarantees that each request should only start once. Equation (6) ensures that at any time $t$, the total memory usage in the KV cache does not exceed the memory limit $M$.
Since solving IP is computationally intractable, we consider relaxing the integer variables $x_{i,t}$ to be continuous variables between 0 and $1$, and we have the following relaxed LP formulation:
$\text{Relax-LP}=\min\,$ $\sum_{i\in[n]}\left(\sum_{t=\{0,\dots,\bar{T}\}}t\cdot x_{i,t}+o_{i}\right)$ (8)
s.t. $\sum_{t=\{0,\dots,\bar{T}\}}x_{i,t}=1,$ $\forall i\in[n]$ (9)
$\sum_{i=1}^{n}\sum_{k=\max\{0,t-o_{i}\}}^{t-1}(s_{i}+t-k)x_{i,k}\leq M,$ $\forall t\in[\bar{T}]$ (10)
$x_{i,t}\geq 0.$ $\forall i\in[n],\forall t\in[\bar{T}]$ (11)
This relaxation preserves all constraints while making the problem computationally tractable. However, the optimal LP solutions $x_{i,t}^{*}$ typically yield fractional values between 0 and 1, requiring specialized techniques to convert them into actionable scheduling decisions for the discrete original problem. To bridge this gap, we introduce Sorted-LP, an LP-based heuristic that systematically transforms fractional LP solutions into executable schedules.
The algorithm proceeds in three key steps: first, it solves the Relax-LP formulation to obtain fractional assignment variables $x_{i,t}^{*}$. These values are interpreted probabilistically, where each $x_{i,t}^{*}$ represents the likelihood that request $i$ initiates processing at time $t$. Second, for each request $i$, we compute its expected starting time as $y_{i}=\sum_{t\in[\bar{T}]}t\cdot x_{i,t}^{*}$. Finally, Sorted-LP prioritizes requests based on these computed $y_{i}$ values. By sorting all requests in ascending order of $y_{i}$, the algorithm naturally favors requests that the relaxed solution recommends starting earlier. This approach provides a principled way to translate optimization insights from the LP relaxation into a practical scheduling policy. The detailed procedure is formalized in Algorithm 6, which can be found in Appendix 11.1).
While Sorted-LP provides a systematic way to derive schedules from LP solutions, we recognize that further refinement may be possible by incorporating our F-metric framework. To this end, we propose LP-Swap, a hybrid approach that enhances LP-based scheduling with F-metric-guided optimizations. Building on the initial ordering from Sorted-LP, LP-Swap applies local swap operations driven by the quality metric $F(\mathcal{X})$. This hybrid strategy leverages both the global perspective captured by the LP solution and the local optimization capabilities of F-metric refinements. By integrating these complementary approaches, LP-Swap aims to achieve schedule quality beyond what standalone LP solutions can deliver. The complete algorithm pseudocode is provided in Algorithm 7 in Appendix 11.1. We evaluate the effectiveness of the proposed LP-based algorithms using simulated data across a range of distributions in Appendix 11.2, and compare them on the actual dataset in Section 7.
7 Numerical Experiments
In this section, we conduct a numerical experiment with open-source real datasets to evaluate the performance of Sorted-F and its approximations, including the newly proposed Sorted-LP and LP-Swap.
Dataset Overview. To evaluate LLM inference performance under varying input prompt lengths—particularly in scenarios mixing short and long prompts—we combine two publicly available datasets, as no single existing dataset meets this need. The first dataset, introduced by Zheng et al. (2023), contains conversational data from over 210,000 unique IP addresses, collected via the Vicuna demo and Chatbot Arena. This dataset, available at https://huggingface.co/datasets/lmsys/lmsys-chat-1m, consists of short, interactive exchanges with a mean input length of 41.84 tokens (median: 12.00) and a mean output length of 85.35 tokens (median: 43.00), as shown in Figure 5.
Figure 5: Distribution of the number of tokens of input prompt and output response respectively in the data Zheng et al. (2023)
The second dataset, constructed by Cohan et al. (2018), focuses on long-form summarization of academic papers from Arxiv (available at https://github.com/armancohan/long-summarization). Here, prompts are significantly longer, with a mean input length of 2546.39 tokens (median: 2667.50) and a mean output length of 296.08 tokens (median: 161.00), illustrated in Figure 6.
Figure 6: Distribution of the number of tokens of input prompt and output response respectively in the data Cohan et al. (2018)
From these datasets, we randomly select 1600 short conversations and 400 long summarization tasks. By merging them with a random permutation, we create a mixed dataset of 2000 samples. Figure 7 reveals the resulting distribution: the input lengths now exhibit a mean of 542.75 tokens (median: 18.00), while outputs have a mean of 127.50 tokens, demonstrating the desired variability in prompt sizes.
Figure 7: Distribution of the number of tokens of input prompt and output response respectively in the mixed data
Experiment Setup. We evaluate performance using averaged end-to-end latency, defined as the total latency across all requests divided by the number of requests. For benchmarking, we compare against three algorithms:
  1. 1.
    FCFS: Requests are processed in the order they arrive (determined by random shuffling of the mixed dataset). Each batch is filled maximally without exceeding memory limits, assuming perfect knowledge of output lengths.
  2. 2.
    MC-SF: As described in Section 2.1, this strategy prioritizes requests with shorter output lengths to minimize batch processing time.
  3. 3.
    Sorted-LP: Introduced in Section 6, Sorted-LP solves a relaxed linear program to estimate request starting times, then batches requests in ascending order of these values.
  4. 4.
    LP-Swap: The hybrid method outlined in Section 6, combining Sorted-LP ordering with F-metric-guided swaps.
To test our proposed algorithm, Sorted-F, we face computational constraints: exact dynamic programming is intractable for 2000 requests. Instead, we employ the local swap search and the quantile greedy selection approximation methods.
For scalability analysis, we evaluate latency trends by subsampling the mixed dataset ($200,400,\ldots,2000$ requests). All simulations emulate deployment of LLaMA2-70B on two A100 GPUs, with a KV cache memory limit of 16,492 tokens. Batch processing times are estimated using the linear regression method in the Vidur simulator from Agrawal et al. (2024a).
Results. Figure 8 compares the averaged latency across scheduling algorithms, revealing key insights.
Figure 8: Averaged latency between different scheduling algorithms
First, the inferior performance of MC-SF relative to Sorted-F and Sorted-LP underscores the importance of accounting for variable input sizes—purely output-length-based prioritization (MC-SF) proves inadequate for mixed workloads.
For our proposed Sorted-F algorithm, the results align with the theoretical analysis in Section 5: the quantile greedy selection achieves faster runtime, while the local swap method delivers superior latency performance. This trade-off highlights the practical value of both approximation approaches.
Notably, LP-Swap demonstrates significant latency reduction over its base algorithm Sorted-LP, validating that incorporating the F-metric substantially improves LP-based schedules. However, the performance of LP-Swap achieves nearly identical latency to Sorted-F (Swap), as demonstrated in the detailed comparison provided in Appendix 12 (Figure 10). This indicates that the initial LP-based ordering provides no unique advantage over direct F-metric optimization, as both converge to similar schedules after swap refinement. This empirically confirms that F-metric-guided optimization is highly effective irrespective of initial scheduling. The robustness of the F-metric as a scheduling criterion is thus reinforced, and Sorted-F remains highly competitive due to its significantly faster computation time and simpler implementation.
Appendix: Proofs of Statements
8 Supplementary Materials for Section 2
Proof 8.1
Proof of Theorem 2.1: Consider the following instance $\mathcal{I}$ with two types of requests:
  • Type 1: each request has size $s=\sqrt{M}-1$, output length $o=1$, and there are $X=M$ such requests.
  • Type 2: each request has size $s=1$, output length $o=2$, and there are $Y=M^{1.5}$ such requests.
Under MC-SF, which processes all type 1 requests before any type 2 request (since type 1 requests are shorter), the total expected latency is
$\text{TEL}(\texttt{MC-SF};\mathcal{I})=\sqrt{M}\sum_{i=1}^{X/\sqrt{M}}i+\frac{XY}{\sqrt{M}}+\frac{2M}{3}\sum_{j=1}^{3Y/M}j.$
Now consider an alternative schedule $\Lambda$ that processes all type 2 requests before any type 1 request. Its total latency is
$\text{TEL}(\Lambda;\mathcal{I})=\frac{2M}{3}\sum_{j=1}^{3Y/M}j+\frac{6XY}{M}+\sqrt{M}\sum_{i=1}^{X/\sqrt{M}}i.$
We compare the two total latencies:
$\frac{\text{TEL}(\texttt{MC-SF};\mathcal{I})}{\text{TEL}(\Lambda;\mathcal{I})}=\frac{\frac{1}{2}M^{1.5}+3M^{1.25}+M^{2}}{\frac{1}{2}M^{1.5}+3M^{1.25}+6M^{1.5}}\geq\frac{2}{13}M^{0.5}.$
[…]
Theorem 8.2
Suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then Algorithm $\texttt{MC-SF}_{2}$ (process the smallest $s_{i}+o_{i}$ first) achieves an unbounded competitive ratio. Specifically, as $M\to\infty$, we have
$\text{CR}(\texttt{MC-SF}_{2})\to\infty.$
Proof 8.3
Proof of Theorem 8.2: Consider the following instance $\mathcal{I}$ with two types of requests:
  • Type 1: each request has size $s=1$, output length $o=\sqrt{M}-1$, and there are $X=M$ such requests.
  • Type 2: each request has size $s=\sqrt{M}$, output length $o=1$, and there are $Y=M^{1.5}$ such requests.
Under $\texttt{MC-SF}_{2}$, which processes all type 1 requests before any type 2 request (since type 1 requests are shorter), the total expected latency is
$\text{TEL}(\texttt{MC-SF}_{2};\mathcal{I})=(\sqrt{M}-1)\sqrt{M}\sum_{i=1}^{X/\sqrt{M}}i+(\sqrt{M}-1)\frac{XY}{\sqrt{M}}+\sqrt{M}\sum_{j=1}^{Y/\sqrt{M}}j.$
Now consider an alternative schedule $\mathcal{A}$ that processes all type 2 requests before any type 1 request. Its total latency is
$\text{TEL}(\mathcal{A};\mathcal{I})=(\sqrt{M}-1)\sqrt{M}\sum_{i=1}^{X/\sqrt{M}}i+\frac{XY}{\sqrt{M}}+\sqrt{M}\sum_{j=1}^{Y/\sqrt{M}}j.$
We compare the two total latencies:
$\frac{\text{TEL}(\texttt{MC-SF}_{2};\mathcal{I})}{\text{TEL}(\mathcal{A};\mathcal{I})}=\frac{\frac{1}{2}M^{2}+M^{2.5}+\frac{1}{2}M^{1.75}}{\frac{1}{2}M^{2}+M^{2}+\frac{1}{2}M^{1.75}}\geq\frac{2}{3}M^{0.5}.$
[…]
Proof 8.4
Proof of Theorem 2.2 We prove this via a reduction from the 3-Partition problem, which is known to be strongly NP-hard. Consider the following 3-Partition Problem: Given a multiset of integers $X=\{x_{1},\dots,x_{3m}\}$ summing to $mT$, where each $x_{i}$ satisfies $T/4<x_{i}<T/2$, the problem asks whether $X$ can be partitioned into $m$ disjoint subsets $S_{1},\dots,S_{m}$ such that the sum of each subset equals $T$.
Then, we construct the following reduction: Given an instance of 3-Partition, we construct an instance as follows:
  • For each integer $x_{i}$, create a request $i$ with $s_{i}=x_{i}$ and $o_{i}=1$.
  • Set the memory capacity $M=T$.
We show that the 3-Partition instance has a solution if and only if the constructed instance has a schedule with total end-to-end latency $\text{TEL}=\frac{3m(m+1)}{2}$.
Case 1: 3-Partition Exists. Suppose $X$ can be partitioned into $m$ subsets $S_{1},\dots,S_{m}$, each summing to $T$. Then, consider the schedule which processes batch $k$ with all requests in $S_{k}$. The schedule is feasible since the memory usage of each batch is $M$. Moreover, as each batch processes exactly 3 requests (due to $T/4<x_{i}<T/2$), the total end-to-end latency is
$\text{TEL}=\sum_{t=1}^{m}3t=\frac{3m(m+1)}{2}.$
Case 2: No 3-Partition Exists. If no such partition exists, then at least one batch must process fewer than 3 requests (since no subset of 2 requests sums to $\leq T$, given $x_{i}>T/4$). This forces the schedule to use at least $m+1$ batches. Compared to the schedule in case 1, there is at least one job having to be swapped from one of the batch $1$ to batch $m$ to the batch $m+1$, and the total end-to-end latency in this case is strictly greater than $\frac{3m(m+1)}{2}$.
[…]
Theorem 8.5
Under the model suggested in Jaillet et al. (2025), suppose that each request $i\in[n]$ satisfies $s_{i}\in[1,M]$, $o_{i}\in[1,M]$ and $s_{i}+o_{i}\leq M$, then minimizing the makespan is NP-hard.
Proof 8.6
Proof of Theorem 8.5 We prove NP-hardness via a reduction from the Partition problem. Consider the following Partition Problem: Given a multiset of integers $X=\{x_{1},\dots,x_{n}\}$ summing to $2T$, does there exist a partition of $X$ into two subsets $S_{1},S_{2}$ such that $\sum_{i\in S_{1}}x_{i}=\sum_{i\in S_{2}}x_{i}=T$?
We construct the following reduction:
  • For each integer $x_{i}$, create a request $i$ with $s_{i}=x_{i}$ and $o_{i}=1$
  • Set memory capacity $M=T$.
We show that the Partition instance has a solution if and only if the constructed instance admits a schedule with makespan 2.
Case 1: Partition Exists. Given subsets $S_{1},S_{2}$ each summing to $T$, the schedule:
  • Batch 1: Process all requests in $S_{1}$ (memory usage = $T$)
  • Batch 2: Process all requests in $S_{2}$ (memory usage = $T$)
completes all requests in makespan 2.
Case 2: No Partition Exists. Any valid schedule must use at least 3 batches because:
  • No single batch can process all requests (total memory $2T>M$)
  • Any two-batch solution would require both batches to have memory exactly $T$, which would constitute a valid partition
[…]
9 Supplementary Materials for Section 4
Proof 9.1
Proof of Lemma 4.7: According to the construction of $\mathcal{Y}_{m}$, for any $k\in\{b_{m-1}+1,b_{m-1}+2,\cdots,b_{m}-1\}$, the following inequality holds:
$\frac{\sum_{j=1}^{n_{k}}o_{k,j}+o_{k+1,1}}{(n_{k}+1)^{2}}>\frac{\sum_{j=1}^{n_{k}}o_{k,j}}{n_{k}^{2}}.$
This inequality implies that incorporating request $r_{k+1,1}$ into batch $\mathcal{X}_{k}$ would not yield improvement in the objective function value $F(\mathcal{X}_{k})$. Consequently, while such addition may remain within the memory constraint $M$, it is excluded by the optimality criteria of the batching procedure.
Let $q_{k+1}$ denote the maximum number of requests from batch $k+1$ that can be added to $\mathcal{X}_{k}$ without exceeding $M$, i.e., $\mathcal{X}_{k}+[r_{k+1,1},\dots,r_{k+1,q_{k+1}}]$ remains feasible. The critical inequality
$\frac{\sum_{i=1}^{n_{k}}o_{k,i}+\sum_{j=1}^{q_{k+1}}o_{k+1,j}}{(n_{k}+q_{k+1})^{2}}>\frac{\sum_{i=1}^{n_{k}}o_{k,i}}{n_{k}^{2}}$
must hold. Otherwise, incorporating these $q_{k+1}$ requests would have improved $\mathcal{X}_{k}$ by either reducing $F(\mathcal{X}_{k})$ or increasing its batch size while maintaining $F(\mathcal{X}_{k})$, which would contradict the optimal selection criterion of Sorted-F.
This inequality directly implies the following relationship between the average processing times:
$\frac{\sum_{j=1}^{q_{k+1}}o_{k+1,j}}{q_{k+1}}>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{\sum_{i=1}^{n_{k}}o_{k,i}}{n_{k}}.$ (12)
Next, we examine the relationship between the longest processing times $o_{k+1,n_{k+1}}$ in batch $\mathcal{X}_{k+1}$ and $o_{k,n_{k}}$ in batch $\mathcal{X}_{k}$. Since $o_{k+1,n_{k+1}}$ is the maximum processing time in $\mathcal{X}_{k+1}$, it dominates both individual terms and their average:
$o_{k+1,n_{k+1}}\geq o_{k+1,q_{k+1}}\geq\frac{1}{q_{k+1}}\sum_{j=1}^{q_{k+1}}o_{k+1,j}.$ (13)
Furthermore, Lemma 3.3 establishes that for large $n_{k}$, the average processing time in $\mathcal{X}_{k}$ satisfies
$\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}o_{k,i}>\frac{1}{2}\cdot o_{k,n_{k}}.$
Combining these results with Inequality (12) yields the key relationship:
$o_{k+1,n_{k+1}}$ $\geq\frac{1}{q_{k+1}}\sum_{j=1}^{q_{k+1}}o_{k+1,j}$
$>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}o_{k,i}$
$>\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{1}{2}o_{k,n_{k}}$
$>o_{k,n_{k}}.$ (14)
Subsequently, we extend the analysis to consecutive batches $\mathcal{X}_{k+1}$ and $\mathcal{X}_{k+2}$ for any $k\in\{b_{m-1}+1,\dots,b_{m}-2\}$. The construction similarly satisfies
$\frac{\sum_{j=1}^{n_{k+1}}o_{k+1,j}+o_{k+2,1}}{(n_{k+1}+1)^{2}}>\frac{\sum_{j=1}^{n_{k+1}}o_{k+1,j}}{n_{k+1}^{2}},$
which demonstrates that including request $r_{k+2,1}$ in $\mathcal{X}_{k+1}$ would not reduce the objective function value $F(\mathcal{X}_{k+1})$. While such inclusion may not exceed memory $M$, it is precluded by the optimality criteria of the batching procedure.
Let $q_{k+2}$ denote the maximum number of requests from $\mathcal{X}_{k+2}$ that can be added to $\mathcal{X}_{k+1}$ without violating $M$. This yields two key inequalities:
$\frac{\sum_{j=1}^{q_{k+2}}o_{k+2,j}}{q_{k+2}}$ $>\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{i=1}^{n_{k+1}}o_{k+1,i}}{n_{k+1}}$
$o_{k+2,n_{k+2}}$ $\geq o_{k+2,q_{k+2}}\geq\frac{\sum_{j=1}^{q_{k+2}}o_{k+2,j}}{q_{k+2}}.$
Furthermore, the increasing response length property implies
$\frac{\sum_{i=1}^{n_{k+1}}o_{k+1,i}}{n_{k+1}}\geq\frac{\sum_{j=1}^{q_{k+1}}o_{k+1,j}}{q_{k+1}}.$
Combining these results with previous Inequalities (12), (13), we can derive the cumulative relationship:
$o_{k+2,n_{k+2}}$ $\geq\frac{\sum_{l=1}^{q_{k+2}}o_{k+2,l}}{q_{k+2}}$
$>\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{i=1}^{n_{k+1}}o_{k+1,i}}{n_{k+1}}$
$\geq\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\frac{\sum_{j=1}^{q_{k+1}}o_{k+1,j}}{q_{k+1}}$
$\geq\left(2+\frac{q_{k+2}}{n_{k+1}}\right)\cdot\left(2+\frac{q_{k+1}}{n_{k}}\right)\cdot\frac{\sum_{i=1}^{n_{k}}o_{k,i}}{n_{k}}$
$>4\cdot\frac{1}{2}o_{k,n_{k}}$
$=2o_{k,n_{k}}.$ (15)
Based on Inequalities 14 and 15, for any $i\in\{1,\dots,a_{m}\}$, we can derive the following geometric bound for all batch indices $i\in\{1,\dots,a_{m}\}$:
$o_{k+i,n_{k+i}}\leq\left(\frac{1}{2}\right)^{\lfloor\frac{a-i}{2}\rfloor}\cdot o_{k+a,n_{k+a}}.$
[…]
10 Supplementary Materials for Section 5
Algorithm 2 Exact Optimal Request Selection via Dynamic Programming
[…]
2:Initialize $dp[k][m]\leftarrow\infty$ and $path[k][m]\leftarrow\emptyset$ for k=0..n,m=0..Mk=0..n,\ m=0..M
3:dp[0][0]0dp[0][$\leftarrow 0$
4:for each request $r_{i}\in\mathcal{I}$ do
5:  $m_{i}\leftarrow s_{i}+o_{i}$
6:  $o_{i}\leftarrow$ output length
7:  for $k\leftarrow n-1$ downto 0 do
8:   for each $m$ where $dp[k][m]<\infty$ do
9:     $new_{m}\leftarrow m+m_{i}$
10:     if $new_{m}\leq M$ then
11:      $new\_val\leftarrow dp[k][m]+o_{i}$
12:      if $new\_val<dp[k+1][new_{m}]$ then
13:        $dp[k+1][new_{m}]\leftarrow new\_val$
14:        $path[k+1][new_{m}]\leftarrow path[k][m]\cup\{i\}$
15:      end if
16:     end if
17:   end for
18:  end for
19:end for
[…]
21:for $k=1$ to $n$ do
22:  for $m=0$ to $M$ do
23:   if $dp[k][m]<\infty$ then
24:     $F\leftarrow dp[k][m]/(k\times k)$
25:     if $F<F^{*}$ then
26:      $F^{*}\leftarrow F$
27:      $\mathcal{X}^{*}\leftarrow\{r_{i}:i\in path[k][m]\}$
28:     end if
29:   end if
30:  end for
31:end for
32:return $\mathcal{X}^{*}$
Algorithm 3 Scaled $(1+\epsilon)$-Approximation via Dynamic Programming
[…]
3:Initialize $dp[k][\hat{m}$$\leftarrow\infty$ and $path[k][\hat{m}$$\leftarrow\emptyset$
4:dp[0][0]0dp[0][$\leftarrow 0$
5:for each request $r_{i}\in\mathcal{I}$ do
6:  $\hat{m}_{i}\leftarrow\lfloor(s_{i}+o_{i})/\lambda\rfloor$
7:  for $k=n$ downto 11 do
8:   for $\hat{m}=\hat{M}$ downto $\hat{m}_{i}$ do
9:     if $dp[k-1][\hat{m}-\hat{m}_{i}]+o_{i}<dp[k][\hat{m}$] then
10:      $dp[k][\hat{m}]\leftarrow dp[k-1][\hat{m}-\hat{m}_{i}]+o_{i}$
11:      $path[k][\hat{m}]\leftarrow path[k-1][\hat{m}-\hat{m}_{i}]\cup\{i\}$
12:     end if
13:   end for
14:  end for
15:end for
[…]
17:for $k=1$ to $n$ do
18:  for $\hat{m}=0$ to $\hat{M}$ do
19:   $F\leftarrow dp[k][\hat{m}]/k^{2}$
20:   if $F<F^{*}$ then
21:     $F^{*}\leftarrow F$
22:     $\mathcal{X}^{*}\leftarrow path[k][\hat{m}$]
23:   end if
24:  end for
25:end for
26:return $\mathcal{X}^{*}$
Algorithm 4 Local Swap Search Heuristic
1:Sort $\mathcal{I}$ by ascending $s_{i}+o_{i}$
[…]
3:for each $r\in\mathcal{I}$ in sorted order do
4:  if $mem+s_{r}+o_{r}\leq M$ then
5:   $\mathcal{X}\leftarrow\mathcal{X}\cup\{r\}$
6:   $mem\leftarrow mem+s_{r}+o_{r}$
7:  end if
8:end for
[…]
10:while improvedimproved do
11:  $improved\leftarrow False$
12:  $currentF\leftarrow(\sum_{r\in\mathcal{X}}o_{r})/|\mathcal{X}|^{2}$
13:  for each $r_{out}\in\mathcal{X}$ do
14:   for each $r_{in}\in\mathcal{I}\setminus\mathcal{X}$ do
15:     $\Delta_{m}\leftarrow(s_{in}+o_{in})-(s_{out}+o_{out})$
16:     if $mem+\Delta_{m}>M$ then
17:      continue
18:     end if
19:     $\Delta_{o}\leftarrow o_{in}-o_{out}$
20:     $newF\leftarrow(\sum_{r\in\mathcal{X}}o_{r}+\Delta_{o})/|\mathcal{X}|^{2}$
21:     if $newF<currentF$ then
22:      $\mathcal{X}\leftarrow(\mathcal{X}\setminus\{r_{out}\})\cup\{r_{in}\}$
23:      $mem\leftarrow mem+\Delta_{m}$
24:      $improved\leftarrow True$
25:      break inner loop
26:     end if
27:   end for
28:   if improvedimproved then
29:     break
30:   end if
31:  end for
32:end while
33:return $\mathcal{X}$
Algorithm 5 Quantile Greedy Selection
1:Sort $\mathcal{I}$ by ascending $o_{i}$
[…]
7:for each $r\in\mathcal{I}$ in sorted order do
8:  if $(s_{r}+o_{r})\leq Q_{p}$ and $o_{r}\leq Q_{o}$ and $mem+s_{r}+o_{r}\leq M$ then
9:   $\mathcal{X}\leftarrow\mathcal{X}\cup\{r\}$
10:   $mem\leftarrow mem+s_{r}+o_{r}$
11:  end if
12:end for
[…]
14:Sort remainingremaining by ascending $o_{i}/(s_{i}+o_{i})$
15:for each $r\in remaining$ in sorted order do
16:  if $mem+s_{r}+o_{r}\leq M$ then
17:   $\mathcal{X}\leftarrow\mathcal{X}\cup\{r\}$
18:   $mem\leftarrow mem+s_{r}+o_{r}$
19:  end if
20:end for
21:return $\mathcal{X}$
11 Supplementary Materials of Section 6
11.1 Pseudocodes for LP-Based Methods
Algorithm 6 Sorted-LP
1:Solve Relax-LP and obtain optimal solutions $x_{i,t}^{*}$ for all $i\in[n],t\in[\bar{T}]$
2:for each $i\in[n]$ do
3:  Compute $y_{i}\leftarrow\sum_{t\in[\bar{T}]}t\cdot x_{i,t}^{*}$
4:end for
5:Sort all requests $r_{i}\in\mathcal{I}$ in ascending order of $y_{i}$ to obtain $\mathcal{I}^{\prime}$
[…]
7:while $\mathcal{R}^{(t)}+\mathcal{V}^{(t)}$ do
8:  $t\leftarrow t+1$, $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}-\mathcal{U}^{(t-1)}$, $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t-1)}+\mathcal{U}^{(t-1)}$, $\mathcal{U}^{(t)}\leftarrow[]$
9:  for $r_{i}\in\mathcal{R}^{(t)}$ in order of $\mathcal{I}^{\prime}$ do
10:   if $M(\mathcal{V}^{(t)}+\mathcal{U}^{(t)},p_{j}+o_{j})\leq M,\forall r_{j}\in\mathcal{V}^{(t)}+\mathcal{U}^{(t)}+[r_{i}]$ then
11:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}+[i]$
12:   else
13:     Break
14:   end if
15:  end for
[…]
17:end while
Algorithm 7 LP-Swap
1:Solve Relax-LP and obtain optimal solutions $x_{i,t}^{*}$ for all $i\in[n],t\in[\bar{T}]$
2:for each $i\in[n]$ do
3:  Compute $y_{i}\leftarrow\sum_{t\in[\bar{T}]}t\cdot x_{i,t}^{*}$
4:end for
5:[$\mathcal{I}^{\prime}\leftarrow$[]
6:while $\mathcal{I}$ do
7:  Sort all requests $r_{i}\in\mathcal{I}$ in ascending order of $y_{i}$
8:  $\mathcal{X}_{\text{init}}\leftarrow\emptyset,\ mem\leftarrow 0$
9:  for each $r\in\mathcal{I}$ in sorted order do
10:   if $mem+s_{r}+o_{r}\leq M$ then
11:     $\mathcal{X}_{\text{init}}\leftarrow\mathcal{X}_{\text{init}}\cup\{r\}$
12:     $mem\leftarrow mem+s_{r}+o_{r}$
13:   end if
14:  end for
15:  $\mathcal{X}^{*}\leftarrow\text{LocalSwap}(\mathcal{X}_{\text{init}},\mathcal{I},M)$
16:  Sort $r_{i}\in\mathcal{X}^{*}$ in ascending order of $o_{i}$
17:  $\mathcal{I}^{\prime}\leftarrow\mathcal{I}^{\prime}+\mathcal{X}$, $\mathcal{I}\leftarrow\mathcal{I}-\mathcal{X}$
18:end while
[…]
20:while $\mathcal{R}^{(t)}+\mathcal{V}^{(t)}$ do
21:  $t\leftarrow t+1$, $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t-1)}-\mathcal{U}^{(t-1)}$, $\mathcal{V}^{(t)}\leftarrow\mathcal{V}^{(t-1)}+\mathcal{U}^{(t-1)}$, $\mathcal{U}^{(t)}\leftarrow[]$
22:  for $r_{i}\in\mathcal{R}^{(t)}$ in sorted order do
23:   if $M(\mathcal{V}^{(t)}+\mathcal{U}^{(t)},p_{j}+o_{j})\leq M,\forall r_{j}\in\mathcal{V}^{(t)}+\mathcal{U}^{(t)}+[r_{i}]$ then
24:     $\mathcal{U}^{(t)}\leftarrow\mathcal{U}^{(t)}+[i]$
25:   else
26:     Break
27:   end if
28:  end for
[…]
30:end while
11.2 Simulation Results and Analysis
This appendix presents a comprehensive analysis of Sorted-LP, Sorted-F (Swap), and LP-Swap algorithms through simulations on synthetically generated data. We employed five distinct distributions—Uniform, Normal, Binomial, Exponential, and Mixed—to model input ($s_{i}$) and output ($o_{i}$) lengths, with detailed parameters as follows.
Experimental Setup and Parameters
For all experiments, memory capacity $M=100$ and maximum sequence lengths $s_{\max}=o_{\max}=50$ were fixed. The distributions used are:
  • Uniform: $s_{i}\sim\mathcal{U}(1,51)$, $o_{i}\sim\mathcal{U}(1,51)$.
  • Normal: $\mu=25$, $\sigma=8.33$, $s_{i},o_{i}\sim\mathcal{N}(\mu,\sigma^{2})$ clipped to $[1,50]$.
  • Binomial: $s_{i},o_{i}\sim\text{Binomial}(49,0.5)+1$.
  • Exponential: $s_{i},o_{i}\sim\text{Exp}(\lambda=0.2)$ (scale=5) clipped to $[1,50]$.
  • Mixed: Designed to mimic real-world patterns in Section 7:
    • Input sequences ($s_{i}$):
        80% exponentially distributed with $\lambda=0.1$ (scale=10).
        20% lognormally distributed with $\mu=\ln 40$, $\sigma=0.25$.
        Values >50>50 remapped to $\mathcal{U}(40,50)$.
    • Output lengths ($o_{i}$): Exponentially distributed with $\lambda=0.2$ (scale=5) clipped to $[1,50]$.
Figure 9 illustrates the distribution of input sequence lengths ($s_{i}$) for the Mixed dataset, highlighting its bimodal structure with predominantly short sequences and occasional longer requests.
Figure 9: Distribution of input sequence lengths ($s_{i}$) in the Mixed dataset configuration
Numerical Results
Table 4 summarizes the average total end-to-end latency (TEL) across all trials, providing the quantitative basis for our subsequent analysis.
Table 4: Average total end-to-end latency across distributions
Distribution Sorted-LP Sorted-F (Swap) LP-Swap Requests ($n$)
Uniform 9763.0 9748.3 9771.4 100
Normal 9397.3 9416.9 9385.3 100
Binomial 10943.0 10696.7 10665.7 100
Exponential 6015.7 6016.9 6016.0 100
Mixed 22133.2 22100.1 22128.2 200
Key Observations
Through detailed analysis of algorithm execution—including batch formation dynamics, request sequencing patterns, and runtime scheduling decisions—we identify three notable empirical patterns:
1. LP’s preference for short-output requests: The Sorted-LP approach consistently prioritizes requests with shorter output lengths ($o_{i}$), even when their input sizes ($s_{i}$) are substantial. This behavior appears related to the LP formulation’s relaxation, where fractional $x_{i,t}$ variables reduce the perceived cost of memory blocking by large inputs. Consequently, minimizing output latency dominates scheduling decisions, which can lead to suboptimal memory allocation when large-input requests are scheduled early.
2. Swap’s balancing effect: The swap mechanism effectively counteracts this bias by considering both input and output sizes during local exchanges. This dual consideration produces more balanced schedules that better account for memory constraints, particularly in distributions where naive prioritization of short outputs could allow memory-intensive requests to cause disproportionate blocking.
3. LP+Swap synergy and performance: The hybrid LP-Swap approach combines LP’s global optimization perspective with swap’s local refinement capabilities. While pure swap achieves marginally better results in the Mixed distribution (designed to simulate real-world data), LP-Swap demonstrates competitive performance across all tested distributions. This consistent adaptability suggests that integrating both techniques offers a promising direction for developing robust schedulers that maintain effectiveness under diverse request patterns.
These empirical patterns provide motivation for further investigating hybrid scheduling strategies, which we pursue with real-world datasets in Section 7.
12 Supplementary Materials of Section 7
Figure 10 is a refined figure of Figure 8, which only compares the two curves (Sorted-F and LP-Swap) which almost overlap with each other.
Figure 10: Averaged latency between Sorted-F and LP-Swap